OpenStudy (anonymous):

The length of a diagonal of a square is 8.5 inches. Find the area of the square. a 36 inches2 b 12 inches2 c 24 inches2 d 90 inches2

OpenStudy (anonymous):

help please

OpenStudy (sugarmochi):

Do you know Pythagoras theorem?

OpenStudy (anonymous):

 L^2=W^2+W^2 pythagorus 8.5^2=2W^2 72.25=2W^2 36= W^2 =area

OpenStudy (anonymous):

thank you very much